OpenStudy (mathrocks2020):

The base length of a triangle is 4 feet and the height is 2 feet. What is the area of the triangle?

OpenStudy (mathstudent55):

Here is the formula for the area of a triangle. Just use your given values in the formula. \(\Large A_{triangle} = \dfrac{base \times height}{2} \)

OpenStudy (mathrocks2020):

So 4×2/2?

OpenStudy (mathrocks2020):

@mathstudent55

OpenStudy (mathrocks2020):

Which ! means the answer is 4?

OpenStudy (calculusxy):

Yes! \(\huge \bf \frac{4 \times 2}{2} = \frac{8}{2} = 4\) \(\huge \bf \color{green}{\checkmark}\)